OpenStudy (anonymous):

Translate the sentence into an equation. Two hundred minus three times a number is equal to nine. -------------------------------------------------------------------------------- 200 - 3n = 9 3n - 200 = 9 3(200 - n) = 9 (200 - 3)n = 9

OpenStudy (anonymous):

I put the first one but idk

OpenStudy (kidrah69):

thats right

OpenStudy (kidrah69):

200 minus= 200- 3 times a number= 3n equals 9 = 9

OpenStudy (anonymous):

so basically it would be two hundred-threexnumber=nine just put the words into number

OpenStudy (anonymous):

good enough for you @DivineSolar ?

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

200 - 3n = 9